During a marathon race, a runner’s blood flow increases to 10.0 times her resting rate. Her blood’s viscosity has dropped to 95.0% of its normal value, and the blood pressure difference across the circulatory system has increased by 50.0%. By what factor has the average radii of her blood vessels increased? 


1
Expert's answer
2021-03-21T11:39:54-0400
"Q=\frac{\Delta P\pi r^4}{8\eta l}\\r_1=\left(\frac{Q_18\eta_1l}{\Delta P_1\pi}\right)^{0.25}
\\r_2=\left(\frac{Q_28\eta_2l}{\Delta P_2\pi}\right)^{0.25}
\\r_2=\left(\frac{10(8)(0.95)}{1.5\pi}\right)^{0.25}r_1\\r_2=1.586r_1"